The difference of two numbers is 33. The second is 7 less than 3 times the first. What are the two numbers?
I think the answers are 20 and 53, but I am getting confused...Can you help?
Thanks

Answer by ankor@dixie-net.com(12705)

That's right, here is a method, easy to understand.
:
"The difference of two numbers is 33".
Let x = 1st number
Then the 2nd number = (x+33)
:
"The second is 7 less than 3 times the first."
(x+33) = 3x - 7
x - 3x = -7 - 33
-2x = -40
x = -40/-2x
x = +20